OpenStudy (anonymous):

Solve, If f'(x) = Cos(x^2) and y = f [ (2x-1) / (x+1) ], find dy/dx

ganeshie8 (ganeshie8):

we may use chain rule \[y=f(\color{blue}{g(x)}) ~\implies ~\dfrac{dy}{dx}=f'(\color{blue}{g(x)})\times \color{blue}{g'(x)}\]

OpenStudy (anonymous):

Okay, I ended up having dy/dx = f' [ (2x-1) / (x+1) ] x [ 3 / (x+1)^2 ]

ganeshie8 (ganeshie8):

looks good, plugin the value of f' [ (2x-1) / (x+1) ]

ganeshie8 (ganeshie8):

f'(x) = cos(x^2) f' [ (2x-1) / (x+1) ] = ?

OpenStudy (anonymous):

Would it equal Cos(x^2)? Or do you have to substitute Cos(x^2) into f'?

OpenStudy (anonymous):

Like as in f' [ (2x-1) / (x+1) ] = Cos(x^2) OR Cos(x^2) [ (2x-1) / (x+1) ]

ganeshie8 (ganeshie8):

not quite, simply replace x by (2x-1) / (x+1)

OpenStudy (anonymous):

Ah okay So like: Cos( [ (2x-1) / (x+1) ] ^2 ) ?

ganeshie8 (ganeshie8):

\[f'(\color{red}{x}) = \cos(\color{red}{x}^2)\] so \[f'(\color{red}{ (2x-1) / (x+1) }) = \cos(\color{red}{ [ (2x-1) / (x+1) ] }^2)\]

OpenStudy (anonymous):

Ah okay thank you ^^